-3k+1+7k=33 i need a step by step solution a check please
Answer by checkley71(8403)

-3k+1+7k=33
4k=33-1
4k=32
k=32/4
k=8 answer.
proof
-3*8+1+7*8=33
-24+1+56=33
-24+57=33
33=33